# Rotatable Antenna Enabled Wireless Communication: Modeling and Optimization

### v4

@@ V-A Single-User System
To compare the directional gains of RA and fixed-antenna systems, Fig. 5 plots the maximum received signal power versus the number of antennas $N_{x}$ for a ULA system. The asymptotic values given in (27) and (28) are also shown in the figure. It is first observed that the closed-form SNR derived in (30) matches perfectly with the exact value calculated by (23), which validates the correctness of Theorem 1 and Lemma 1. Additionally, for a small to moderate number of antennas, the received signal powers of both the RA and fixed-antenna systems increase linearly with $N_{x}$, which is in accordance with Remark 2. However, as $N_{x}$ further increases, it is observed that the received signal powers of both RA and fixed-antenna systems eventually approach their asymptotic values. Meanwhile, the RA system reaches its asymptotic limit later and achieves up to 1.43 dB gain over the fixed-antenna system, which corroborates the accuracy of analytical result in (29) as $10\mathrm{log}_{10}\frac{\lim_{N_{x}\to\infty}\tilde{\gamma}}{\lim_{N_{x}\to\infty}\tilde{\gamma}_{\mathrm{fixed}}}=10\mathrm{log}_{10}\left(\frac{\pi}{6}+\operatorname{cos}\frac{\pi}{6}\right)\approx$ 1.43 dB. Furthermore, when $N_{x}\leq 100$, the RA system achieves up to 5 dB gain over the conventional fixed-antenna system. This indicates that the proposed RA architecture can achieve a high array gain due to the additional spatial DoFs induced by the antenna boresight rotation.
@@
In Fig. 6, we plot the received signal power versus the antenna directivity factor $p$ for the UPA system. It is observed that the received signal power of the RA system increases as $p$ increases. This is expected since a higher directivity (i.e., larger $p$) results in greater radiation power in the boresight direction, leading to a higher array gain focused on the user direction when the boresight of each antenna is aligned as closely as possible with the user direction. Conversely, the received signal power of the fixed-antenna system first increases and then decreases with $p$. This is because, with a relatively low antenna directivity (e.g., $p\leq 4$), the main lobe is sufficiently broad to encompass the user direction, allowing the fixed antenna to offer enhanced directional gain compared to the isotropic antenna. However, as $p$ continues to increase, the main lobe becomes narrower, eventually excluding the user direction from its effective coverage. Consequently, the directional gain in the user direction decreases with the further increase of $p$. Moreover, it is observed that the received signal power of the RA system is consistently higher than that of the fixed-antenna system regardless of $p$, highlighting the performance advantage of the RA system, which benefits from the radiation power focusing capability enabled by orientation/boresight adjustments.
@@
Fig. 7 shows the received signal power versus the azimuth angle of the user $\phi$ for the UPA system, where the zenith angle of the user is given by $\psi=\frac{\pi}{2}$. As $\phi$ increases from 0 to $\frac{\pi}{2}$ or decreases from 0 to $-\frac{\pi}{2}$, the received signal power of the fixed-antenna system drastically decreases. This is due to the fact that the array directional gain pattern of the fixed-antenna system is fixed and the radiation power only focuses on the region directly in front of the array. In contrast, the RA array enables the BS to maintain more stable and uniform received signal power over the entire angular region, particularly when $\phi$ lies within the interval $[-\frac{\pi}{6},\frac{\pi}{6}]$, where each RA can adjust its boresight within the allowable rotational range to precisely align with the user direction. In addition, the RA system achieves much higher received signal power than the fixed-antenna system even when the user direction deviates significantly from the array’s main direction, i.e., $\phi\to-\frac{\pi}{2}$ or $\frac{\pi}{2}$. These advantages stem from the RA array’s ability to flexibly reconfigure its directional gain pattern to enhance the directional gain in the user direction. The above results indicate that RA array has the potential to uniformly enhance communication coverage performance in its front half-space.
V-B Multi-User System
Next, we consider a multi-user system under the multipath channel model with $K=4$ users and $Q=8$ scatterer clusters. Specifically, four users are uniformly distributed in four distinct directions in front of the BS. The distance of each user to the BS is independently and randomly selected from a uniform distribution within $[30,50]$ m. Around these users, eight scatterer clusters are randomly distributed. In the following, we present simulation results by averaging over 500 independent channel realizations. Meanwhile, the max-min effective communication rate (which monotonically increases with the SINR) among all the $K$ users is considered as the performance metric, which is given by
$C$ $=\min_{k}\;(1-\varrho)\mathrm{log}_{2}\left(1+\gamma_{k}\right)$
$=(1-\varrho)\mathrm{log}_{2}\left(1+\min_{k}\;\gamma_{k}\right),$ (55)
where the weighting coefficient $\varrho$ with $10\leq\varrho<$ 1 is introduced to account for the system overhead. In our simulations, the weighting coefficients for the RA and fixed-antenna systems are set as $\varrho_{\mathrm{RA}}=0.05$ and $\varrho_{\mathrm{fixed}}=0$, respectively.
In Fig. 8, we illustrate the convergence behavior of the proposed AO algorithm (i.e., Algorithm 1) with different numbers of RAs. It is observed that the max-min effective communication rate increases over iterations and converges within six iterations. This indicates that the proposed AO algorithm converges quickly and demonstrates its effectiveness. To further validate the performance advantages of our proposed RA system, we consider the following four benchmark schemes for comparison:
@@ V-B Multi-User System
  • Array-wise orientation adjustment: In this scheme, we adjust the orientation of the entire antenna array instead of that of each antenna element. By applying the MMSE receive beamforming, the optimal orientation of the antenna array is obtained by using exhaustive search.
@@
  • Baseline with isotropic antennas: In this scheme, the antenna elements in the array are isotropic, i.e., $p=0$ and the radiation energy is evenly distributed in the front half-space of the antennas, and the MMSE receive beamforming is applied at the BS.
@@
In Fig. 9, we compare the max-min effective communication rates versus the user transmit power $P$ for different optimization algorithms. First, it is observed that the AO algorithm with MMSE receive beamforming obtains the highest max-min effective communication rate and achieves up to 2.5 dB gain over the two-stage algorithm. As discussed in Remark 3, this gain gap is expected since the lack of iteration and the assumption of $p=1$ in the two-stage algorithm inevitably cause performance loss. Second, by optimizing the pointing vectors of all RAs to maximize the minimum weighted channel power gain with much lower computational complexity than the AO algorithm, the two-stage algorithm still achieves up to 5 dB gain over the fixed-antenna system without orientation/boresight optimization. The above results indicate the different trade-offs between performance and complexity offered by the two proposed algorithms. Specifically, the AO algorithm can achieve better performance due to its iterative optimization process, while the two-stage algorithm offers competitive performance at significantly lower computational overhead. Furthermore, since the ZF receiver enhances the noise power, the AO algorithm with MMSE beamforming always outperforms that with ZF beamforming.
@@
Figure 9: Max-min effective communication rates obtained by different optimization algorithms versus the user transmit power $P$.
@@
Fig. 10 shows the max-min effective communication rates of different schemes versus the maximum zenith angle $\theta_{\mathrm{max}}$. Several interesting observations are made as follows. First, as $\theta_{\mathrm{max}}$ increases, the proposed RA system gains more DoFs and flexibilities to balance the array directional gain over the multipath channels, thus leading to a further increase in its max-min effective communication rate. Second, the proposed RA system always outperforms both the array-wise orientation adjustment counterpart and the fixed-antenna system. This is because neither of the latter two systems can independently adjust the orientation/boresight of each individual antenna to reconfigure the directional gain pattern of the entire array. Third, since the RAs with random orientations can statistically radiate power in any direction of the BS to serve spatially-distributed users, it can achieve a higher max-min effective communication rate than the fixed-antenna system. However, when $\theta_{\mathrm{max}}\geq\frac{3\pi}{10}$, the max-min effective communication rate of the random orientation design scheme declines with $\theta_{\mathrm{max}}$. This result highlights the importance of antenna orientation/boresight optimization in an RA system, since the random orientations will lead to an unordered array directional gain pattern and inevitable performance loss when $\theta_{\mathrm{max}}$ becomes large. Last but not least, it is interesting to note that the growth rate of the max-min effective communication rate obtained by our proposed RA system substantially increases when $\theta_{\mathrm{max}}\leq\frac{\pi}{10}$, which indicates that even with a small rotational range for RA orientation/boresight adjustment, the proposed RA system with optimized pointing directions can achieve significant performance improvement.
@@
Figure 10: Max-min effective communication rates of different schemes versus the maximum zenith angle $\theta_{\mathrm{max}}$.
@@
Figure 11: Max-min effective communication rates of different systems versus the antenna directivity factor $p$.
@@
Fig. 11 shows the max-min effective communication rates of different schemes versus the antenna directivity factor $p$. It is observed that the max-min effective communication rate of the proposed RA system increases with the directivity factor $p$. This is attributed to the fact that a larger $p$ leads to a higher directional gain in the antenna’s boresight direction and a narrower main lobe, which is more advantageous for the proposed RA system. Consequently, the RA system with larger $p$ can more effectively enhance directional gains across multiple user directions, thereby achieving a larger max-min effective communication rate. In contrast, the max-min effective communication rates of the array-wise orientation adjustment scheme and the fixed-antenna system decrease with $p$ when $p\geq 1$. The reason for this is that with a larger directivity factor $p$, the radiation power of both schemes will be more concentrated in the region directly in front of the array. As a result, the directional gains for users deviating from the main direction of the array will become weaker, thus resulting in a lower max-min effective communication rate. Additionally, although the random orientation design scheme can disperse the radiation power of the array in multiple specific directions, it is significantly inferior to the proposed RA system since it fails to strategically allocate the antenna resources to fairly improve the communication performance of all users. The above results highlight the necessity of our proposed RA system for increasing channel capacity, especially in situations where the antennas have strong directivity, i.e., their main lobes are narrow.
@@ VI Conclusion
In this paper, we proposed a new RA model that provides new spatial DoFs to reconfigure the array directional gain pattern by flexibly adjusting the 3D orientation/boresight of each antenna, thus substantially enhancing the array gain and transmission rate without increasing the number of antennas or changing antenna positions. Specifically, the receive beamforming and the pointing vectors of all RAs were jointly optimized to maximize the minimum SINR among all users. The optimal closed-form pointing vectors for all RAs were first derived with MRC receive beamforming applied in the single-user and free-space propagation setup. Meanwhile, the asymptotic performance analysis for the case with an infinite number of antennas demonstrated that the RA system consistently achieves a higher array gain compared to the conventional fixed-antenna system. For the general multi-user and multipath channel setup, an AO algorithm and a two-stage algorithm were proposed to obtain high-quality suboptimal solutions to balance the array directional gain over the multipath channels. Simulation results validated our analytical results and demonstrated that our proposed RA system significantly outperforms various benchmark schemes. It was shown that even with a small rotational range for RA orientation/boresight adjustment, the RA system could still reap considerable performance gains.
@@ Appendix A Proof of Theorem 1
$\mathbb{D}=$ $2\left[\int_{0}^{\frac{\bar{N}_{x}\Delta}{2}}{\frac{1}{1+\frac{x^{2}}{r^{2}}}\mathrm{d}x}+\int_{\frac{\bar{N}_{x}\Delta}{2}}^{\frac{N_{x}\Delta}{2}}{\frac{\operatorname{cos}\left(\operatorname{arctan}\left(|\frac{x}{r}|\right)-\theta_{\mathrm{max}}\right)}{1+\frac{x^{2}}{r^{2}}}\mathrm{d}x}\right]$
=$\overset{(a)}{=}$ $2r\left[\int_{0}^{\frac{\bar{N}_{x}\delta}{2}}{\frac{1}{1+x^{2}}\mathrm{d}x}+\int_{\frac{\bar{N}_{x}\delta}{2}}^{\frac{N_{x}\delta}{2}}{\frac{\operatorname{cos}\theta_{\mathrm{max}}+\operatorname{sin}\theta_{\mathrm{max}}x}{(1+x^{2})^{\frac{3}{2}}}\mathrm{d}x}\right]$
=$\overset{(b)}{=}$ $2r\left[\left.\operatorname{arctan}\left(x\right)\right|_{0}^{\frac{\bar{N}_{x}\delta}{2}}+\left.\frac{\operatorname{cos}\theta_{\mathrm{max}}x-\operatorname{sin}\theta_{\mathrm{max}}}{\sqrt{1+x^{2}}}\right|_{\frac{\bar{N}_{x}\delta}{2}}^{\frac{N_{x}\delta}{2}}\right]$
=$\overset{(c)}{=}$ $2r\left[\theta_{\mathrm{max}}+\operatorname{sin}\left(\operatorname{arctan}\left(\frac{N_{x}\delta}{2}\right)-\theta_{\mathrm{max}}\right)\right],$ (57)
where $(a)$ holds due to $\operatorname{cos}(\operatorname{arctan}(x))=\frac{1}{\sqrt{1+x^{2}}}$ and $\operatorname{sin}(\operatorname{arctan}(x))=\frac{x}{\sqrt{1+x^{2}}}$, $(b)$ follows the integral formulas 2.103.4, 2.264.5, and 2.264.6 in [47], and $(c)$ holds due to the fact that $\operatorname{arctan}\left(\frac{\bar{N}_{x}\delta}{2}\right)=\theta_{\mathrm{max}}$.
Thus, based on (A) and (A), the proof of Theorem 1 is completed.

### v5

@@ V-A Single-User System
Fig. 5 plots the maximum received signal power versus the number of antennas $N_{x}$ for a ULA system. The asymptotic values given in (27) and (28) are also shown in the figure. It is first observed that the closed-form SNR derived in (30) matches perfectly with the exact value calculated by (23), which validates the correctness of Theorem 1 and Lemma 1. Additionally, for a small to moderate number of antennas, the received signal powers of both the RA and fixed-antenna systems increase linearly with $N_{x}$, which is in accordance with Remark 1. However, as $N_{x}$ further increases, it is observed that the received signal powers of both RA and fixed-antenna systems eventually approach their asymptotic values. Meanwhile, the RA system reaches its asymptotic limit later and achieves up to 1.43 dB gain over the fixed-antenna system, which corroborates the accuracy of analytical result in (29) as $10\mathrm{log}_{10}\frac{\lim_{N_{x}\to\infty}\tilde{\gamma}}{\lim_{N_{x}\to\infty}\tilde{\gamma}_{\mathrm{fixed}}}=10\mathrm{log}_{10}\left(\frac{\pi}{6}+\operatorname{cos}\frac{\pi}{6}\right)\approx$ 1.43 dB. Furthermore, when $N_{x}\leq 100$, the RA system achieves up to 5 dB gain over the conventional fixed-antenna system.
@@
In Fig. 6, we plot the received signal power versus the antenna directivity factor $p$ for the UPA system. It is observed that the received signal power of the RA system increases as $p$ increases. This is expected since a higher directivity (i.e., larger $p$) results in greater radiation power in the boresight direction, leading to a higher array gain focused on the user direction. Conversely, the received signal power of the fixed-antenna system first increases and then decreases with $p$. This is expected since with a relatively low antenna directivity (e.g., $p\leq 4$), the main lobe is sufficiently broad to encompass the user direction, allowing the fixed antenna to offer enhanced directional gain compared to the isotropic antenna. However, as $p$ continues to increase, the main lobe becomes narrower, eventually excluding the user direction from its effective coverage. Moreover, it is observed that the received signal power of the RA system is consistently higher than that of the fixed-antenna system regardless of $p$. This highlights the performance advantage of the RA system, which benefits from the radiation energy focusing capability.
@@
Fig. 7 shows the received signal power versus the azimuth angle of the user $\phi$ for the UPA system, where the zenith angle of the user is given by $\psi=\frac{\pi}{2}$. As $\phi$ increases from 0 to $\frac{\pi}{2}$ or decreases from 0 to $\frac{\pi}{2}$, the received signal power of the fixed-antenna system drastically decreases. This is due to the fact that the array directional gain pattern of the fixed-antenna system is fixed and the radiation power only focuses on the region directly in front of the array. In contrast, the RA array enables the BS to maintain more stable and uniform received signal power over the entire angular region. These advantages stem from the RA array’s ability to flexibly reconfigure its directional gain pattern to enhance the directional gain in the user direction.
V-B Multi-User Narrow-Band System
Next, we consider a multi-user system under the multipath channel model with $K=4$ users and $Q=8$ scatterer clusters. Specifically, four users are uniformly distributed in four distinct directions in front of the BS. The distance of each user to the BS is independently and randomly selected from a uniform distribution within $[30,50]$ m. Around these users, eight scatterer clusters are randomly distributed. In the following, we present simulation results by averaging over 500 independent channel realizations. Meanwhile, we aim to maximize the rate achievable by all $K$ users as the performance metric:
$C=\min_{k}\;\mathrm{log}_{2}\left(1+\gamma_{k}\right)=\mathrm{log}_{2}\left(1+\min_{k}\;\gamma_{k}\right).$ (55)
In Fig. 8, we show the convergence behavior of the proposed AO algorithm (i.e., Algorithm 1) with different numbers of RAs. It is observed that the max-min achievable rate increases over iterations and converges within six iterations. This indicates that the proposed AO algorithm converges quickly and is effective. To further validate the performance advantages of our proposed RA system, we consider the following three benchmark schemes for comparison:
@@ V-B Multi-User Narrow-Band System
  • Baseline with isotropic antennas: In this scheme, the antennas in the array are isotropic, i.e., $p=0$ and the radiation energy is evenly distributed in the front half-space of the antennas, and the MMSE receive beamforming is applied at the BS.
@@
In Fig. 9, we compare the max-min achievable rates versus the user transmit power $P$ for different optimization algorithms. First, it is observed that the AO algorithm with MMSE receive beamforming achieves the highest max-min achievable rate and achieves up to 2.5 dB gain over the two-stage algorithm. As discussed in Remark 2, this gain gap is expected since the lack of iteration and the assumption of $p=1$ in the two-stage algorithm inevitably cause performance loss. Second, by optimizing the RA pointing vectors to maximize the weighted channel power gain, the two-stage algorithm still achieves up to 5 dB gain over the fixed-antenna system. The above results indicate the different trade-offs between performance and complexity offered by the two proposed algorithms. Specifically, the AO algorithm can achieve better performance due to its iterative optimization process, while the two-stage algorithm offers competitive performance at significantly lower computational overhead. Third, to achieve the same max-min achievable rate, the RA system requires significantly lower user transmit power than both the fixed-antenna and isotropic-antenna systems. This is because antenna rotation enables the BS to dynamically adjust its directional gain based on user distribution, thereby capturing more signal power from the users.
@@
Figure 9: Max-min achievable rates obtained by different optimization algorithms versus the user transmit power $P$.
@@
Fig. 10 shows the max-min achievable rates of different schemes versus the maximum zenith angle $\theta_{\mathrm{max}}$. Several interesting observations are made as follows. First, as $\theta_{\mathrm{max}}$ increases, the proposed RA system gains more DoFs and flexibilities to balance the array directional gain over the multipath channels, thus leading to a further increase in its max-min achievable rate. Second, by adjusting antenna orientations/boresights to reconfigure the overall directional gain pattern, the proposed RA system consistently outperforms the fixed-antenna system. Third, since the RAs with random orientations can statistically radiate power in any direction of the BS to serve spatially-distributed users, it can achieve a higher max-min achievable rate than the fixed-antenna system. However, when $\theta_{\mathrm{max}}\geq\frac{3\pi}{10}$, the max-min achievable rate of the random orientation design declines with $\theta_{\mathrm{max}}$. This result highlights the importance of antenna orientation/boresight optimization in an RA system, since the random orientations will lead to an unordered array directional gain pattern and inevitable performance loss when $\theta_{\mathrm{max}}$ becomes large. Last but not least, the growth rate of the max–min communication rate in our proposed RA system rises sharply when $\theta_{\mathrm{max}}\leq\frac{\pi}{10}$, showing that even a small rotational range for RA orientation/boresight adjustment can yield substantial performance gains.
@@
Figure 10: Max-min achievable rates of different schemes versus the maximum zenith angle $\theta_{\mathrm{max}}$.
@@
Figure 11: Max-min achievable rates of different systems versus the antenna directivity factor $p$.
@@
Fig. 11 shows the max-min achievable rates of different schemes versus the antenna directivity factor $p$. It is observed that the max-min achievable rate of the proposed RA system increases with the directivity factor $p$. This is attributed to the fact that a larger $p$ leads to a higher directional gain in the antenna’s boresight direction and a narrower main lobe, which is more advantageous for the proposed RA system. Consequently, the RA system with larger $p$ can more effectively enhance directional gains across multiple user directions, thereby achieving a larger max-min effective communication rate. In contrast, the max-min achievable rate of the fixed-antenna system decreases with $p$ when $p\geq 1$, as a larger directivity factor concentrates the radiated power more narrowly in front of the array. As a result, the directional gains for users deviating from the main direction of the array will become weaker, thus resulting in a lower max-min achievable rate. Additionally, although the random orientation design can disperse the radiation power of the array in multiple directions, it is significantly inferior to the proposed RA system since it fails to strategically allocate the antenna resources to fairly improve the communication performance of all users.
V-C Polarization-Aware Wideband Channel
To validate the effectiveness and performance advantages of the RA architecture in practical scenarios, we extend the multi-user narrowband setting to the polarization-aware wideband channel setup, as described in Appendix E. Specifically, we consider $K=4$ users and $Q=8$ scatterer clusters, with their geometric distributions similarly given in Section V-B. We adopt orthogonal frequency division multiplexing (OFDM) with the carrier frequency $f_{c}=2.4$ GHz and bandwidth $B=40$ MHz. The number of subcarriers and the cyclic prefix (CP) length are configured as $L=64$ and $L_{\mathrm{CP}}=6$, respectively. Each user has a transmit power budget of $P_{k}=10$ dBm, and the maximum zenith angle of each RA is $\theta_{\mathrm{max}}=\frac{\pi}{6}$. Vertically polarized antennas are employed at both the BS and users. The achievable sum-rate defined in (77) is adopted as the performance metric. For comparison, we consider the benchmark schemes described in Section V-B, applied with the same MRC beamforming and subcarrier allocation as described in Appendix E. In the polarization-unaware RA system, both the RA pointing vectors and subcarrier allocation are optimized based on the polarization-unaware channel models defined in (8) and (9).
Figure 12: Achievable sum-rates of different schemes versus the antenna directivity factor $p$.
Fig. 12 shows the achievable sum-rate versus the antenna directivity factor, demonstrating the performance of RA systems under polarization-aware modeling. It can be observed that the achievable sum-rates of RA systems exhibit a unimodal dependence on the directivity factor $p$. Specifically, the performance improves as $p$ increases up to an intermediate value (around $p=5$), but then degrades. This behavior reflects the fundamental trade-off between radiation gain and spatial coverage: insufficient directivity fails to provide adequate gain toward the users, while excessive directivity limits the ability to capture multiple propagation paths in wideband communication. Notably, the polarization-aware RA system consistently outperforms all other schemes across the entire range of $p$. This advantage stems from the flexible adjustment of antenna orientation/boresight direction, which effectively balances polarization alignment and directional gain to improve channel quality. These results confirm the effectiveness of the proposed RA architecture in practical wideband systems with polarization effects.
Figure 13: Achievable sum-rates of different schemes versus the number of the subcarriers $L$.
Fig. 13 illustrates the achievable sum-rate versus the number of subcarriers, validating the robustness of the proposed RA design under wideband OFDM settings. It is observed that the achievable sum-rates of all schemes increase with the number of subcarriers. This improvement stems from higher spectral efficiency, as the relative overhead of the CP decreases with more subcarriers. Additionally, for RA systems, increasing $L$ improves subcarrier allocation flexibility, which can be jointly optimized with antenna orientations. The polarization-aware RA system consistently achieves the best performance by jointly optimizing pointing vectors and subcarrier allocation while accounting for polarization variation due to antenna rotation. Although the polarization-unaware RA system also outperforms fixed and random orientation designs as well as the isotropic baseline, its sum-rate remains slightly lower than that of the polarization-aware counterpart. This highlights the importance of accurate polarization modeling when optimizing RA configurations. Nevertheless, the polarization-unaware design still provides a computationally efficient alternative with acceptable performance.
@@ VI Conclusion
In this paper, we proposed a new RA model that provides new spatial DoFs by flexibly adjusting the 3D orientation/boresight of each antenna. Specifically, the receive beamforming and the pointing vectors of all RAs were jointly optimized to maximize the minimum SINR among all users. The optimal closed-form pointing vectors for all RAs were first derived in the single-user and free-space propagation setup. For the general multi-user and multipath channel setup, an AO algorithm and a two-stage algorithm were proposed to obtain high-quality suboptimal solutions to balance the array directional gain over the multipath channels. Simulation results validated our analytical results and demonstrated that our proposed RA system significantly outperforms various benchmark schemes. It was shown that even with a small rotational range for RA orientation/boresight adjustment, the RA system could still reap considerable performance gains.
Building upon the general channel model and theoretical performance of the proposed RA system, future research may delve into developing new practical techniques to enhance system performance. Promising directions include the design of more efficient channel estimation algorithms and the low-overhead boresight rotation schemes, as well as the extension of the current framework to emerging 6G applications.
@@ Appendix A Proof of Theorem 1
$\hskip-5.69046pt\mathbb{D}=$ $2\hskip-1.42271pt\left[\int_{0}^{\frac{\bar{N}_{x}\Delta}{2}}{\frac{1}{1+\frac{x^{2}}{r^{2}}}\mathrm{d}x}\hskip-1.42271pt+\hskip-2.84544pt\int_{\frac{\bar{N}_{x}\Delta}{2}}^{\frac{N_{x}\Delta}{2}}{\frac{\operatorname{cos}\left(\operatorname{arctan}\left(|\frac{x}{r}|\right)\hskip-1.42271pt-\hskip-1.42271pt\theta_{\mathrm{max}}\right)}{1+\frac{x^{2}}{r^{2}}}\mathrm{d}x}\right]$
$\overset{(a)}{=}$ $2r\left[\int_{0}^{\frac{\bar{N}_{x}\delta}{2}}{\frac{1}{1+x^{2}}\mathrm{d}x}+\int_{\frac{\bar{N}_{x}\delta}{2}}^{\frac{N_{x}\delta}{2}}{\frac{\operatorname{cos}\theta_{\mathrm{max}}+\operatorname{sin}\theta_{\mathrm{max}}x}{(1+x^{2})^{\frac{3}{2}}}\mathrm{d}x}\right]$
$\overset{(b)}{=}$ $2r\left[\left.\operatorname{arctan}\left(x\right)\right|_{0}^{\frac{\bar{N}_{x}\delta}{2}}+\left.\frac{\operatorname{cos}\theta_{\mathrm{max}}x-\operatorname{sin}\theta_{\mathrm{max}}}{\sqrt{1+x^{2}}}\right|_{\frac{\bar{N}_{x}\delta}{2}}^{\frac{N_{x}\delta}{2}}\right]$
$\overset{(c)}{=}$ $2r\left[\theta_{\mathrm{max}}+\operatorname{sin}\left(\operatorname{arctan}\left(\frac{N_{x}\delta}{2}\right)-\theta_{\mathrm{max}}\right)\right],$ (57)
where $(a)$ holds due to $\operatorname{cos}(\operatorname{arctan}(x))=\frac{1}{\sqrt{1+x^{2}}}$ and $\operatorname{sin}(\operatorname{arctan}(x))=\frac{x}{\sqrt{1+x^{2}}}$, $(b)$ follows the integral formulas 2.103.4, 2.264.5, and 2.264.6 in [10], and $(c)$ holds due to the fact that $\operatorname{arctan}\left(\frac{\bar{N}_{x}\delta}{2}\right)=\theta_{\mathrm{max}}$. Thus, based on (A) and (A), the proof of Theorem 1 is completed.